a particle moves with constant speed v along a regular hexagon ABCDEF  in same order . then magnitude of average velocity for its motion from A to C IS ?
 

Grade:12th pass

3 Answers

Sourabh Singh IIT Patna
7 years ago
Hii
The displacwmwnt between the given points is not 3*side but it is sqrt(3)*side.
ence the average velocity between A to C is going to be V/sqrt(3) .


Anshika Maheshwari
6 years ago
When a particle travels it will come to its original place and so it's average velocity will become zero due to zero displacement 
 
But if a particle has to travel from A to C then for average velocity we will find resultant. 
AC=(AB2+BC2+2.AB.BC.cos 60)1/2 
AC=(v2+v2+2v2*1/2)½
AC=v(3)1/2 
Vaverage=(Vfinal —  Vinital)/2
=(V(3)1/2—0)/2
=V(3)1/2/2
yash
5 years ago
bcwbecbiucvbnidcmniabviuerbcvsdnuicnuvbniunciunvbuyciubcsbvysbusdbvyuerbfvbnuicbuixbbodbnvjsnuivnrnuvnsdjivnuevnuenvunvuenvuenunuenvuenvuenvuenvuevnunvuevnunvuenvunvuenvuevnuevnuv